# Do Language Models Encode Collaborative Signals in Recommendation? Findings and Potentials

### v2

@@ C.1 Datasets
We incorporate six datasets in this paper, including four datasets from the Amazon platform 333www.amazon.com (Ni et al., 2019) (i.e., Books, Movies & TV, Video Games, and Industrial & Scientific), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing). Table 6 reports the dataset statistics.
@@ C.2 Baselines
We incorporate a series of ID-based CF models as our baselines for general recommendation. These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec). For these LM-enhanced CF methods, we adopt the leading method XSimGCL as the backbone.
@@ C.3 Ablation Study
We conduct the same ablation study as introduced in Section 4 on Movies & TV and Video Games datasets. As illustrated in Figure 6, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4.
C.4 The t-SNE Visualization Comparison
@@ D.2.2 Baselines
Since previous works about zero-shot recommendation mostly focus on sequential recommendation (Kang & McAuley, 2018; Wang et al., 2019a), we slightly modify two methods in sequential recommendation, ZESRec (Ding et al., 2021) and UniSRec (Hou et al., 2022) as our baselines. Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.

### v3

@@ B.2 Generating Item Representations from LMs
B.3 Connection with Fine-tuned LLM-based Approaches for Exploring the Relationship Between Language Space and Behavior Space.
In exploring the space relationship between language modeling and user behavior modeling, there is another research line that uses LLMs to interpret behavior representations (e.g., item representations trained from MF (Koren et al., 2009)) from traditional recommenders (Tennenholtz et al., 2024; Yang et al., 2023; Lei et al., 2024). They tend to project behavior representations into the token space of LLMs, and fine-tune LLMs to generate explanations for behavior representations (e.g., generate a movie description given a movie representation trained from MF). The fact that LLMs can interpret such behavior representations indicates that the behavior space and language space can be aligned at the token space level, which is similar to the findings in our paper. Our paper and these studies can be viewed as complementary works on a similar research goal (i.e., exploration on the connection between language modeling and user behavior modeling from the representation space perspective), providing support for each other.
@@ C.1 Datasets
We incorporate six datasets in this paper, including four datasets from the Amazon platform 222www.amazon.com (Ni et al., 2019) (i.e., Books, Movies & TV, Video Games, and Industrial & Scientific), and two datasets from other platforms (i.e., MovieLens-1M and Book Crossing). Table 6 reports the dataset statistics.
@@ C.2 Baselines
We incorporate a series of ID-based CF models as our baselines for general recommendation. These models are classified as classical CF methods (MF, MultVAE, and LightGCN), CL-based CF methods (SGL, BC Loss, and XSimGCL), and LM-enhanced CF methods (KAR, RLMRec). We do not consider baselines using LMs as recommenders for two practical reasons: the huge inference cost on datasets with millions of interactions and the task limitation of candidate selection (Liao et al., 2024) or next item prediction (Zheng et al., 2023). For these LM-enhanced CF methods, we adopt the leading method XSimGCL as the backbone. We use the same item and user representations as we adopt in AlphaRec for KAR and XSimGCL, to make the comparison relatively fair. To evaluate the impact of the modifications, we demonstrate one ablation study on RLMRec in Appendix C.5.
@@ C.3 Ablation Study
We conduct the same ablation study as introduced in Section 4 on Movies & TV and Video Games datasets. As illustrated in Figure 6, each component in AlphaRec contributes positively, which is consistent with our findings in Section 4. Specifically, the performance degradation caused by replacing the MLP with a linear weight matrix (w/o MLP) indicates that nonlinear transformations can extract the implicit user preference similarities encoded in the language representation space more effectively. Besides, the performance also drops from replacing InfoNCE loss (Wu et al., 2022) with BPR loss (Rendle et al., 2012) (w/o CL) and removing the graph convolution (w/o GCN) suggests that explicitly modeling the collaborative relationships through the loss function and model architecture can further enhance recommendation performance.
C.4 Effect of Varying Language Representations in AlphaRec
We evaluate the effect of different language representations in AlphaRec by varying adopted LMs. As reported in Table 7, advanced language representations consistently yield higher performance than early BERT-style language representations. Moreover, due to the non-linearity and neighborhood aggregation introduced in AlphaRec, the performance gap between different language representations is narrowed.
Table 7: Effect of varying language representations in AlphaRec.
Movies & TV Video Games Books
Recall NDCG HR Recall NDCG HR Recall NDCG HR
AlphaRec BERT 0.0994 0.0923 0.4873 0.0960 0.0550 0.2179 0.0719 0.0607 0.3434
RoBERTa 0.0967 0.0895 0.4793 0.0947 0.0545 0.2167 0.0710 0.0596 0.3386
Llama2-7B 0.1160 0.1092 0.5388 0.1395 0.0817 0.3003 0.0940 0.0793 0.4081
Mistral-7B 0.1161 0.1097 0.5421 0.1413 0.0828 0.3020 0.0945 0.0799 0.4090
text-embedding-ada-v2 0.1152 0.1083 0.5382 0.1437 0.0844 0.3062 0.0933 0.0784 0.4061
text-embeddings-3-large 0.1221 0.1144 0.5587 0.1519 0.0894 0.3207 0.0991 0.0828 0.4185
SFR-Embedding-Mistral 0.1225 0.1139 0.5571 0.1521 0.0887 0.3209 0.0982 0.0820 0.4161
C.5 Ablation Study on the Modification of RLMRec
We have modified RLMRec (Ren et al., 2024b) by using the same item and user representations as in AlphaRec to make the comparison more fair, rather than using representations of LLM-generated item and user profiles. To evaluate the effectiveness of this modification, we conduct an ablation study on the adopted item and user representations. Specifically, we implement another version of RLMRec (Ren et al., 2024b) by using the original prompting approach to generate item and user profile, and name this version as RLMRec (LLM-generated profile). As shown in Table 8, the performance of RLMRec (LLM-generated profile) is lower than the implementation in this paper. Furthermore, the performance of incorporating LLM-generated profiles is not always stable, which may even lead to performance degradation in some cases (e.g., Movies & TV). We attribute the performance gap to the possible noise and hallucination introduced by these generated profiles.
Table 8: Performance comparison of different versions of RLMRec .
Movies & TV Video Games Books
Recall NDCG HR Recall NDCG HR Recall NDCG HR
XSimGCL 0.1057 0.0984 0.5128 0.1138 0.0662 0.2550 0.0879 0.0745 0.3918
RLMRec (LLM-generated profile) 0.1046 0.0942 0.5063 0.1218 0.0696 0.2692 0.0905 0.0741 0.4049
RLMRec (This paper) 0.1119 0.1013 0.5301 0.1384 0.0809 0.2997 0.0928 0.0774 0.4092
C.6 Performance on Amazon Electronic 2023
To alleviate the information leakage problem that the early datasets may be used in the pretraining data of LMs, we consider the latest version of Amazon 2023 (Hou et al., 2024a) as the newly published dataset. This dataset was published in March 2024, which is the latest public dataset that we have access to. To further alleviate the information leakage problem, we only consider the interactions after 2022. We conduct the same experiment as in Section 4 on this new dataset, and present the performance in Table 9. As shown in this table, in this new dataset, AlphaRec still consistently outperforms baselines. We attribute this to the fact that language models understand the user preference behind item title description, rather than naively remembering the training data. Therefore, AlphaRec still works fine on the latest dataset.
Table 9: Performance of AlphaRec on Amazon Electronics 2023
Amazon Electronics 2023
Recall NDCG HR
MF (Rendle et al., 2012) $0.0130$ $0.0089$ $0.0136$
MultVAE (Liang et al., 2018) $0.0227$ $0.0158$ $0.0237$
LightGCN (He et al., 2021) $0.0237$ $0.0161$ $0.0248$
SGL (Wu et al., 2021) 0.0519 0.0250 0.0551
BC Loss (Zhang et al., 2022) 0.0548 0.0265 0.0585
XSimGCL (Yu et al., 2024) 0.0534 0.0261 0.0569
KAR (Xi et al., 2023) 0.0611 0.0283 0.0661
RLMRec (Ren et al., 2024b) 0.0633 0.0288 0.0674
AlphaRec 0.0687* 0.0323* 0.0732*
C.7 The t-SNE Visualization Comparison
@@ D.2.2 Baselines
Since previous works about zero-shot recommendation mostly focus on sequential recommendation (Kang & McAuley, 2018; Wang et al., 2019a), we slightly modify three methods in sequential recommendation, ZESRec (Ding et al., 2021), UniSRec (Hou et al., 2022), and VQ-Rec (Hou et al., 2023) as our baselines. Specifically, we maintain the model structure as provided in the paper, and adopt the training paradigm of CF.
@@
  • VQ-Rec (Hou et al., 2023) learns vector-quantized representations for items, enabling transferring between datasets. A codebook is learned from language representations, for looking up item embeddings.
@@ Appendix E Hyperparameter Settings and Implementation Details
Table 16: Cost for extracting language representations on Amazon Movie & TV
per 10,000 items Time Cost Money Cost
Llama2-7B 5 mins 0 \$
text-embedding-3-large 40 s 0.17 \$